# Measuring Gravitational Wave Spectrum from Electroweak Phase Transition and Higgs Self-Couplings

### v2

@@ 3.2 Astrophysical Foreground and Background
Compact binary coalescences (CBCs), both within our Galaxy and beyond, produce numerous signals with low SNRs that overlap incoherently. Their superposition gives rise to two main components in the frequency range of space-based GW detectors: a foreground from unresolved galactic binaries, and a background from extragalactic CBCs Boileau:2021sni.
In the simplified case, the contribution from the astrophysical background can be captured by a power-law model Boileau:2020rpg:
@@
where $f_{\text{ref}}=25\ \text{Hz}$ is the reference frequency, and $\varepsilon$ is the spectral index. For a background generated by the incoherent superposition of many binary black hole and neutron star mergers, one typically expects that $\varepsilon=\tfrac{2}{3}$. In our simulations, we inject such a background with an amplitude $\Omega_{\text{ast}}=10^{-8}$, chosen within the log-uniform prior range $(10^{-10},10^{-7})$, consistent with the parameter space considered in Refs. LIGOScientific:2019vic; Boileau:2022ter.
On the other hand, the foreground is mainly due to unresolved galactic white dwarf binaries, which dominate the low-frequency band and form a confusion signal Barack:2004wc; Adams:2013qma; Liu:2023qap. For simplicity, in this work we include only the extragalactic astrophysical background and neglect the galactic foreground.
@@ 4 Detection of SGWB with Space-based Detectors
Here we present the method used for simulating and inferring the SGWB. Space-based GW detectors—such as Taiji, LISA and TianQin—are designed as triangular constellations of three spacecraft connected by laser links Gowling:2022pzb, with each arm length corresponding to the distance between two free-falling test masses Gowling:2021gcy. A passing GW induces tiny differential changes in these arm lengths, which modify the interference pattern recorded by the spacecraft.
The detection of SGWB can be performed either with a single detector or a network of detectors Ruan:2020smc; Wang:2021njt. In this work, we focus on the single-detector case, which differs conceptually from the standard method used by ground-based detectors such as LIGO/Virgo/KAGRA KAGRA:2021kbb, where cross-correlation between multiple detectors is employed to suppress uncorrelated noises Allen:1997ad; Romano:2016dpx. For a space-based constellation, the noises in the different data channels are correlated and cross-correlation is not applicable; instead, a null channel Adams:2010vc can be constructed that is insensitive to GWs and thus provides a probe of instrumental noise, while the remaining channels retain GW sensitivity and can be used to search for the SGWB. We adopt this null-channel method in our analysis.
@@ 4.1 Signal Response and Noise Models
Here we provide a brief introduction to the signal response and noise models for space-based GW detectors, and refer readers to Ref. Smith:2019wny for more detailed discussions of this topic. The GWs are characterized by their amplitude $h_{ij}$, or more appropriately by the spectral density $\Omega_{\text{GW}}(f)$ for SGWB. The detector measures the signals as three streams of data $d_{i}(t)$, where $i=1,2,3$ represent $A_{BC}$, $B_{CA}$, and $C_{AB}$ respectively, and $A,B,C$ label the spacecraft forming the triangular interferometric constellation. These data streams capture the relative optical path change of a Michelson interferometer consisting of two adjacent arms. Due to the properties of the SGWB, the signal amplitudes $d_{1/2/3}(t)$ are random variables. We will work mainly in the frequency domain, with the following convention for the Fourier transform:
@@
The signal part is related to the GW spectrum $\Omega_{\text{GW}}(f)$ by
$S_{a}(f)=\frac{3H_{0}^{2}}{4\pi^{2}}\frac{\Omega_{\text{GW}}(f)}{f^{3}}\mathcal{R}_{a}(f),$ (4.5)
through the response function $\mathcal{R}_{a}(f)$, which, in the idealized configuration we adopt here, takes the following analytical forms Smith:2019wny; Cornish:2001bb
$\mathcal{R}_{A}(f)$ $=\mathcal{R}_{E}(f)=\frac{9}{20}|W(f)|^{2}\left[1+\left(\frac{f}{4f_{\ast}/3}\right)^{2}\right]^{-1},$ (4.6)
$\mathcal{R}_{T}(f)$ $=\frac{1}{4032}|W(f)|^{2}\left(\frac{f}{f_{\ast}}\right)^{6}\left[1+\frac{5}{16128}\left(\frac{f}{f_{\ast}}\right)^{8}\right]^{-1}.$
Here, the factor $W(f)$ in the results above arises from the inherent complexities of space-based GW detection. The primary source of detector noise stems from the fluctuation of the laser frequency combined with the unequal lengths of the interferometer arms. It should be suppressed by approximately eight orders of magnitude through the use of TDI Tinto:2004wu, a method that generates complex combinations of raw signals to achieve noise cancelation. For simplicity, we assume equal and fixed arm lengths in this analysis. The application of TDI introduces the factor $W=1-e^{-2if/f_{\ast}}$, where $f_{\ast}=\tfrac{c}{2\pi L}$, $c$ denotes the speed of light, and $L$ represents the arm length of the detectors, which is $3\times 10^{9}\ \text{m}$ for Taiji and $2.5\times 10^{9}\ \text{m}$ for LISA. The response functions for the Taiji and LISA detectors are presented in Fig. 1. Notably, at low frequencies, the $T$-channel exhibits significantly lower sensitivity to GW signals compared to the $A$ and $E$ channels, thus serving as the null-channel for GW signals.
@@
With the signal contribution now determined, we need the PSD of each of the major noise components $N_{a}(f)$. For both Taiji and LISA, the noise models share a similar form (see e.g., Babak:2021mhe) and consist of two primary components. One is the acceleration noise, given by
@@ 4.2 Data Simulation and Signal Injection
Data-cleaning and pre-processing steps are required during the actual operation of the space-based detectors, such as removing data gaps from interruptions of detector running, etc., which result in a set of time segments $N_{0}$, each with a duration $T$. These segments collectively span a total observation time $T_{t}=N_{0}T$. Here we choose simply $T=10^{6}\ \text{s}$ (approximately 11.4 days) following Caprini:2024hue, and $N_{0}=126$, so that the total effective observing duration is about 4 years.
To accurately reconstruct signals and avoid distortion, the sampling frequency $f_{s}$ must satisfy $f_{s}>2f_{\text{max}}$, where $f_{\text{max}}$ is the maximum frequency of interest, as dictated by the Nyquist sampling theorem. The data are represented as an equally spaced time series with a time interval of $\Delta t$ between samples, resulting in a sampling rate $f_{s}=\tfrac{1}{\Delta t}$. In Fourier transformation analysis, the data from a single observation period are integrated, either over the range 0 to $T$ or from $\tfrac{T}{2}$ to $\tfrac{T}{2}$, to transform into the frequency domain. The minimal frequency resolution achievable in this domain is $\Delta f=\tfrac{1}{T}$, and the maximum analyzable frequency is then $f_{\text{max}}=\tfrac{1}{2\Delta t}$, where $N=\tfrac{T}{\Delta t}$ is the total number of data points within the observation period. In this research, we select $\Delta t=1\ \text{s}$ for convenience, and the frequency range is $\left[3\times 10^{-5},0.5\right]\ \text{Hz}$.
@@
where $f_{k}=\tfrac{k}{N\Delta t}$ and $t_{\mathrm{n}}=n\Delta t$ represent the discrete frequency and time points, respectively. With this convention, the relation between $\tilde{d}(f)$ and $\tilde{d}(f_{k})$ becomes $\tilde{d}(f)=\tilde{d}(f_{k})\,\Delta t$. Thus, Eq. (4.2) can be reformulated as Pieroni:2020rob:
@@
$\left\langle\tilde{d_{a}}(f_{k})\tilde{d}^{*}_{a}(f_{k})\right\rangle=\frac{Tf_{s}^{2}}{2}P_{a}\left(f_{k}\right),\qquad(a=A,E,T).$ (4.13)
Figure 2: This figure shows the theoretical PSD curve (blue) and the simulated data in channels $A$ and $T$ (red points). The curve is computed using Eqs. (4.4) and (4.5), while the red points are generated from Eq. (4.16). The analysis focuses on the frequency range $\left[3\times 10^{-5}\,\text{Hz},\,0.5\,\text{Hz}\right]$, with $\Omega_{\text{ast}}=10^{-8}$ and $\varepsilon=2/3$.
For $\tilde{d}(f)$, which is a complex quantity, its real and imaginary parts are statistically independent and follow the same Gaussian distribution Maggiore:2018sht, with zero mean and variances given by:
$\left\langle\left(\text{Re}\penalty 10000\ \tilde{d}_{a}(f_{k})\right)^{2}\right\rangle=\left\langle\left(\text{Im}\penalty 10000\ \tilde{d}_{a}(f_{k})\right)^{2}\right\rangle=\frac{Tf_{s}^{2}}{4}P_{a}(f_{k})\equiv\sigma_{a}^{2}(f_{k}).$ (4.14)

### v3

@@ 3.2 Astrophysical Foreground and Background
CBCs, both within our Galaxy and beyond, produce numerous signals with low SNRs that overlap incoherently. Their superposition gives rise to two main components in the frequency range of space-based GW detectors: a foreground from unresolved galactic binaries, and a background from extragalactic CBCs Boileau et al. (2021b).
In the simplified case, the contribution from the astrophysical background can be captured by a power-law model Boileau et al. (2021a):
@@
where $f_{\text{ref}}=1\ \text{mHz}$ is the reference frequency, and $\varepsilon$ is the spectral index. For a background generated by the incoherent superposition of many binary black hole and neutron star mergers, one typically expects that $\varepsilon=\tfrac{2}{3}$. In our simulations, we inject such a background with an amplitude $\Omega_{\text{ast}}\simeq 3.16\times 10^{-12}$, chosen within the log-uniform prior range $(10^{-14},10^{-10})$, consistent with the parameter space considered in Refs. Abbott and others (2019); Boileau et al. (2023).
On the other hand, the foreground is mainly due to unresolved galactic white dwarf binaries, which dominate the low-frequency band and form a confusion signal Barack and Cutler (2004); Adams and Cornish (2014); Liu et al. (2023). For simplicity, in this work we include only the extragalactic astrophysical background and neglect the galactic foreground.
@@ 4 Detection of SGWB with Space-based Detectors
Here we present the method used for simulating and inferring the SGWB. Space-based GW detectors—such as Taiji, LISA and TianQin—are designed as triangular constellations of three spacecraft connected by laser links Gowling et al. (2023), with each arm length corresponding to the distance between two free-falling test masses Gowling and Hindmarsh (2021). A passing GW induces tiny differential changes in these arm lengths, which modify the interference pattern recorded by the spacecraft.
The detection of SGWB can be performed either with a single detector or a network of detectors Ruan et al. (2020b); Wang and Han (2021). In this work, we focus on the single-detector case, which differs conceptually from the standard method used by ground-based detectors such as LIGO/Virgo/KAGRA Abbott and others (2021), where cross-correlation between multiple detectors is employed to suppress uncorrelated noises Allen and Romano (1999); Romano and Cornish (2017). For a space-based constellation, the noises in the different data channels are correlated and cross-correlation is not applicable; instead, a null channel Adams and Cornish (2010) can be constructed that is insensitive to GWs and thus provides a probe of instrumental noise, while the remaining channels retain GW sensitivity and can be used to search for the SGWB. We adopt this null-channel method in our analysis.
@@ 4.1 Signal Response and Noise Models
Here we provide a brief introduction to the signal response and noise models for space-based GW detectors, and refer readers to Ref. Smith et al. (2019) for more detailed discussions of this topic. The GWs are characterized by their amplitude $h_{ij}$, or more appropriately by the spectral density $\Omega_{\text{GW}}(f)$ for SGWB. The detector measures the signals as three streams of data $d_{i}(t)$, where $i=1,2,3$ represent $A_{BC}$, $B_{CA}$, and $C_{AB}$ respectively, and $A,B,C$ label the spacecraft forming the triangular interferometric constellation. These data streams capture the relative optical path change of a Michelson interferometer consisting of two adjacent arms. Due to the properties of the SGWB, the signal amplitudes $d_{1/2/3}(t)$ are random variables. We will work mainly in the frequency domain, with the following convention for the Fourier transform:
@@
The signal contribution in channel $a$ is related to the GW energy density spectrum $\Omega_{\mathrm{GW}}(f)$ through
$S_{a}(f)=\frac{3H_{0}^{2}}{4\pi^{2}}\frac{\Omega_{\mathrm{GW}}(f)}{f^{3}}\mathcal{R}_{a}(f),$ (4.5)
where $\mathcal{R}_{a}(f)$ is the detector response function.
In the idealized equal-arm configuration adopted in this work, the analytical expressions for the response functions read Smith et al. (2019); Cornish (2002)
$\mathcal{R}_{A}(f)$ $=\mathcal{R}_{E}(f)=\frac{9}{5}|W(f)|^{2}\left[1+\left(\frac{f}{4f_{\ast}/3}\right)^{2}\right]^{-1},$ (4.6)
$\mathcal{R}_{T}(f)$ $=\frac{1}{1008}|W(f)|^{2}\left(\frac{f}{f_{\ast}}\right)^{6}\left[1+\frac{5}{16128}\left(\frac{f}{f_{\ast}}\right)^{8}\right]^{-1}.$
Here, we adopt the convention $\Delta L/L$, where $\Delta L$ is the difference between the optical path lengths of the two interferometer arms and $L$ is the single-arm length. This differs from the convention $\Delta L/L_{\delta}$ used in Ref. Smith et al. (2019), where $L_{\delta}$ denotes the round-trip light travel distance of the two paths in the Michelson configuration. As a result, our response functions differ by an overall factor of $1/4$ compared to those in Ref. Smith et al. (2019). Both conventions are physically equivalent, provided that the same normalization is used consistently for the response functions and the noise PSD.
The primary source of detector noise arises from fluctuations in the laser frequency combined with unequal interferometer arm lengths. This noise component is suppressed by approximately eight orders of magnitude through the use of TDI Tinto and Dhurandhar (2005), which constructs appropriate linear combinations of raw measurements to achieve laser noise cancelation. For simplicity, we assume equal and fixed arm lengths in this analysis. The application of TDI introduces the factor $W=1-e^{-2if/f_{\ast}}$, where $f_{\ast}=\tfrac{c}{2\pi L}$, $c$ denotes the speed of light, and $L$ represents the arm length of the detectors, which is $3\times 10^{9}\ \text{m}$ for Taiji and $2.5\times 10^{9}\ \text{m}$ for LISA. The response functions for the Taiji and LISA detectors are presented in Fig. 1. At low frequencies, the $T$ channel exhibits a significantly weaker response to GW signals compared to the $A$ and $E$ channels, and therefore serves as an effective null channel for GW detection.
@@
With the signal contribution now determined, we need the PSD of each of the major noise components $N_{a}(f)$. For both Taiji and LISA, the noise models share a similar form (see e.g., Babak et al. (2021)) and consist of two primary components. One is the acceleration noise, given by
@@ 4.2 Data Simulation and Signal Injection
Data-cleaning and pre-processing steps are required during the actual operation of space-based detectors, such as removing data gaps from interruptions in detector operation, which result in a set of time segments $N_{0}$, each with a duration $T$. These segments collectively span a total observation time $T_{t}=N_{0}T$. Here, we choose simply $T=10^{6}\penalty 10000\ \mathrm{s}$ (approximately 11.4 days) following Caprini et al. (2024) and $N_{0}=126$, so that the total effective observing duration is about 4 years. This segmentation approach is also convenient for frequency-domain analyzes since each segment can be assumed to be approximately stationary.
To accurately reconstruct signals and avoid distortion, the sampling frequency $f_{s}$ must satisfy $f_{s}>2f_{\text{max}}$, where $f_{\text{max}}$ is the maximum frequency of interest, as dictated by the Nyquist sampling theorem. The data are represented as an equally spaced time series with a time interval of $\Delta t$ between samples, resulting in a sampling rate of $f_{s}=\tfrac{1}{\Delta t}$. In Fourier transformation analysis, the data from a single observation period are integrated, either over the range 0 to $T$ or from $\tfrac{T}{2}$ to $\tfrac{T}{2}$, to transform into the frequency domain. The minimal frequency resolution achievable in this domain is $\Delta f=\tfrac{1}{T}$, and the maximum analyzable frequency is then $f_{\text{max}}=\tfrac{1}{2\Delta t}$, where $N=\tfrac{T}{\Delta t}$ is the total number of data points within the observation period. In this research, we select $\Delta t=1\ \text{s}$ for convenience, and the frequency range is $\left[3\times 10^{-5},0.5\right]\ \text{Hz}$.
@@
where $f_{k}=\tfrac{k}{N\Delta t}$ and $t_{\mathrm{n}}=n\Delta t$ represent the discrete frequency and time points, respectively. With this convention, the relation between $\tilde{d}(f)$ and $\tilde{d}(f_{k})$ becomes $\tilde{d}(f)=\tilde{d}(f_{k})\,\Delta t$. Thus, Eq. (4.2) can be reformulated as Pieroni and Barausse (2020):
@@
$\left\langle\tilde{d_{a}}(f_{k})\tilde{d}^{*}_{a}(f_{k})\right\rangle=\frac{Tf_{s}^{2}}{2}P_{a}\left(f_{k}\right),\,(a=A,E,T).$ (4.13)
Figure 2: This figure shows the theoretical PSD curve (blue) and the simulated data in channels $A$ and $T$ (red points). The curve is computed using Eqs. (4.4) and (4.5), while the red points are generated from Eq. (4.16). The analysis focuses on the frequency range $\left[3\times 10^{-5}\,\text{Hz},\,0.5\,\text{Hz}\right]$, with $\Omega_{\text{ast}}=10^{-11.5}$ and $\varepsilon=2/3$.
For $\tilde{d}(f)$, which is a complex quantity, its real and imaginary parts are statistically independent and follow the same Gaussian distribution Maggiore (2018), with zero mean and variances given by:
$\begin{split}\left\langle\left(\text{Re}\penalty 10000\ \tilde{d}_{a}(f_{k})\right)^{2}\right\rangle&=\left\langle\left(\text{Im}\penalty 10000\ \tilde{d}_{a}(f_{k})\right)^{2}\right\rangle\\ &=\frac{Tf_{s}^{2}}{4}P_{a}(f_{k})\equiv\sigma_{a}^{2}(f_{k}).\end{split}$ (4.14)